# Bi-AQUA: Bilateral Control-Based Imitation Learning for Underwater Robot Arms via Lighting-Aware Action Chunking with Transformers

### v1

@@ Abstract
Underwater robotic manipulation is fundamentally challenged by extreme lighting variations, color distortion, and reduced visibility. We introduce Bi-AQUA, the first underwater bilateral control-based imitation learning framework that integrates lighting-aware visual processing for underwater robot arms. Bi-AQUA employs a hierarchical three-level lighting adaptation mechanism: a Lighting Encoder that extracts lighting representations from RGB images without manual annotation and is implicitly supervised by the imitation objective, FiLM modulation of visual backbone features for adaptive, lighting-aware feature extraction, and an explicit lighting token added to the transformer encoder input for task-aware conditioning. Experiments on a real-world underwater pick-and-place task under diverse static and dynamic lighting conditions show that Bi-AQUA achieves robust performance and substantially outperforms a bilateral baseline without lighting modeling. Ablation studies further confirm that all three lighting-aware components are critical. This work bridges terrestrial bilateral control-based imitation learning and underwater manipulation, enabling force-sensitive autonomous operation in challenging marine environments. For additional material, please check: https://mertcookimg.github.io/bi-aqua
@@ I Introduction
Underwater robotic manipulation is uniquely challenging because the visual appearance of the same scene can change drastically within seconds under shifts in the spectrum, intensity, and direction of underwater lighting [1]. Such variations—caused by wavelength-dependent attenuation, scattering, turbidity, and bubbles—break the visual consistency required by standard visuomotor policies, often causing catastrophic action drift even on simple pick-and-place tasks. Existing underwater image enhancement methods improve image quality [2, 3], but they do not address the fundamental control problem: closed-loop manipulation policies must adapt both perception and action generation to rapidly varying lighting.
Traditional underwater manipulation is dominated by teleoperation [4, 5], motivating recent work on AUVs, underwater robot arms, and visuomotor learning [6, 7, 8]. However, existing approaches either rely solely on vision or use unilateral control without force feedback, and none provide mechanisms to explicitly model lighting as a latent factor influencing visual observations.
@@
Figure 2: Unilateral Control-based Imitation Learning
@@
II Related Work
II-A Underwater Vision and Robot
Underwater perception is degraded by wavelength-dependent attenuation, scattering, turbidity, and non-uniform lighting [1], motivating extensive work in image enhancement: physics-based [21], learning-based [22], hybrid approaches [23, 24, 3], large datasets [25], and low-light or non-uniform illumination correction [2, 26]. While some methods target robotics [3, 2], they optimize image quality rather than manipulation performance.
Underwater manipulation research has traditionally relied on teleoperation [4, 27, 5], prompting efforts toward autonomous control of underwater vehicles and manipulators [28, 29]. Recent works explore RL and IL for navigation and manipulation [30, 6, 31] and propose systems such as AquaBot [7] or simulated benchmarks [8]. However, as summarized in the Table I, existing systems either depend solely on vision or lack integration of force feedback and lighting modeling, leaving a gap in real underwater manipulation under dynamic lighting.
Our Bi-AQUA treats lighting as a latent variable inside the visuomotor policy, learning a label-free lighting representation that conditions perception (FiLM) and sequence-level action generation (lighting token) under dynamic underwater illumination.
@@ II-B Bilateral Control-based Imitation Learning
Unilateral teleoperation such as ALOHA and Mobile ALOHA [10, 11, 12] relies solely on visual feedback, and the lack of force feedback limits robustness in contact-rich manipulation.
Bilateral control-based imitation learning (Bi-IL) alleviates this limitation by exchanging both position and force between leader and follower robots, enabling demonstrations rich in interaction dynamics [36, 19, 9]. However, most Bi-IL systems use little or no visual input and instead depend almost entirely on joint and force signals. Recent work has begun integrating images into Bi-IL, with Bi-ACT extending ACT—originally developed for unilateral teleoperation [10, 11, 12]—into a Bi-IL framework [14, 15]. These methods combine force-rich demonstrations with high-dimensional perception but have only been evaluated in air with stable lighting.
Underwater environments introduce scattering, color attenuation, turbidity, and dynamic lighting, severely degrading visual observations. Consequently, proprioception-only Bi-IL becomes perceptually blind, and existing vision-based Bi-IL lacks mechanisms to explicitly model lighting, despite its major role in underwater image formation. To our knowledge, no prior work unifies Bi-IL, underwater deployment, and explicit lighting modeling within a single visuomotor policy. Bi-AQUA is the first Bi-IL framework explicitly designed for real underwater manipulation under dynamic lighting.
@@ II-C Robust Visual Policies and Lighting Modeling
Lighting variability is one of the dominant factors degrading visuomotor policy performance, particularly in underwater environments where color attenuation, scattering, and non-uniform illumination cause rapid, appearance-altering shifts in RGB observations. Prior work has attempted to improve robustness through domain randomization [38, 39, 40, 41] and physically based relighting [32]. While effective for data augmentation, these strategies do not provide the policy with an explicit, internal representation of lighting—leaving the visuomotor pipeline to implicitly cope with color and illumination changes.
FiLM [42] offers a flexible mechanism for conditional feature modulation and has been widely used in preference-aware and multi-task manipulation [43, 44, 45]. However, existing applications typically treat lighting only implicitly—through global conditioning or augmentation—without identifying lighting as a distinct latent factor that affects both perception and downstream control.
In contrast, we explicitly model lighting as a structured internal variable. A label-free Lighting Encoder extracts lighting embeddings directly from raw RGB and is trained solely through the imitation objective, requiring no lighting labels. These embeddings modulate the visual backbone via FiLM and are injected as a dedicated lighting token into the transformer encoder for sequence-level conditioning.
Together, these components allow the visuomotor policy to adapt its perception and control to rapidly changing underwater illumination, in ways that existing domain-randomized or augmentation-only systems cannot. Bi-AQUA is therefore the first bilateral control-based imitation learning framework to incorporate explicit, hierarchical lighting modeling for real underwater manipulation.
@@ III-B Data Collection
where $\theta$ and $\tau$ denote joint position and torques, respectively. Position is obtained from encoders, velocities by differentiation, and torques by a disturbance observer (DOB) [46] and a reaction torque observer (RTOB) [47], without dedicated force sensors.
@@ IV-A1 Hardware Configuration
We evaluate Bi-AQUA on a bilateral leader–follower system, as shown in Fig. 8. Each arm has three joints (J1–J3) and a gripper (J4), driven by Dynamixel XW540-T260 (IP68) servo motors. The follower arm is mounted on a 200 mm pedestal inside a water tank of $900\times 450\times 450$ $m$.
@@
A multi-color RGBW LED aquarium light (FEDOUR RD10-400RGBW, 38 cm, 7 W, 21 LEDs, IP68) is installed beneath the follower pedestal and provides eight reproducible lighting conditions. All lighting modes used in our experiments (red, blue, green, white, rgbw, cyan, purple, changing) correspond to the manufacturer’s 100% brightness presets. For dynamic changing mode, we use the device’s built-in “M5” mode, which automatically cycles through all colors at 2 s intervals. This setup ensures consistent spectral output across trials while allowing us to systematically evaluate robustness under both static and rapidly changing underwater lighting.
@@
Bi-AQUA receives RGB observations from two cameras(TIER IV C1 120): a gripper-mounted camera for close-up views and a top camera capturing the global workspace, including the basket.
IV-A2 Task and Evaluation Protocol
Figure 9: Teleoperation data collection for the pick-and-place task under five lighting modes (e.g., white).
@@ IV-A2 Task and Evaluation Protocol
The task is a pick-and-place designed to stress lighting robustness. The follower must grasp a block on the right side of the tank and place it into a basket on the opposite side, passing through four stages: (#1) initial pose, (#2) pick up, (#3) transport, (#4) place.
For teleoperated data collection, as shown in Fig. 9, we execute the water task twice under each of five lighting conditions (red, blue, green, white, rgbw), yielding 10 demonstrations. The primary training object is a polyurethane block, and the operator can adapt motions to visibility while bilateral control provides consistent force feedback.
For autonomous evaluation, we test all eight lighting conditions, including the dynamic changing mode and two unseen colors (cyan, purple). To assess appearance generalization, we additionally evaluate: (a) a black rubber block, (b) a blue sponge, and (c) bubbles around the polyurethane block as shown in Fig. 10. Each configuration is evaluated over 5 autonomous rollouts and considered successful if the object ends inside the basket without being dropped.
Figure 10: Example task for generalization to novel objects and disturbances
Figure 11: Overview of collected datasets
IV-B Dataset and Training Configuration
IV-B1 Demonstration Collection and Preprocessing
We collect 10 bilateral teleoperation episodes (two per training lighting condition). All demonstrations execute the same pick-and-place task; the operator follows an identical intended motion pattern across episodes so that the primary variation comes from lighting rather than from changes in the manipulation strategy. The policy is trained and evaluated solely from follower-side observations; leader-side joint states serve as action labels.
@@ IV-B1 Demonstration Collection and Preprocessing
Two RGB cameras record at 30 Hz with $1920\times 1280$ resolution: (1) a top camera above the water tank, (2) a gripper camera near the follower gripper. Images are resized to $480\times 320$. Joint position, velocity, and torque for $N=4$ joints (J1–J4) are recorded at 1,000 Hz from both arms. Leader joint states are used as actions; follower joint states are part of the observation. For training, joint data are downsampled to 100 Hz, and images are temporally aligned to 100 Hz by selecting the frame closest to each 10 ms interval. Each time step in the final dataset therefore contains (a) two RGB views (top and gripper), (b) follower joint position, velocity, and torque, and the corresponding leader joint states. Across the 10 demonstrations, this results in 15,402 time steps, corresponding to 30,804 RGB frames in total (two views per time step) and 15,402 synchronized joint‑state vectors; each contains 24 values (leader+follower, 4 joints × position/velocity/torque), totaling 369,648 values. The overview of collected datasets is shown in Fig. 11
@@
Although each episode is collected under a known lighting mode (e.g., red, blue, green), we do not provide any lighting IDs or manual labels to the policy. The Lighting Encoder is trained end-to-end from the control objective and learns a latent representation of lighting purely from visual input.
IV-B2 Network Architecture
TABLE II: Ablation of lighting-aware components in Bi-AQUA.
@@ IV-B2 Network Architecture
All model variants share the same Bi-ACT-style visuomotor backbone; the ablations in Table II differ only in the presence or absence of three lighting-aware components. Concretely, the architecture is controlled by three binary switches: (a) a Lighting Encoder that produces a 64-D lighting embedding, (b) FiLM modulation applied to the visual backbone, and (c) a lighting token added to the transformer encoder input. Table II summarizes which of these components is enabled for each variant.
IV-B3 Training Configuration
All models are trained offline using the same 10 demonstrations, preprocessing pipeline, and optimization hyperparameters to ensure a fair comparison across the rows of Table II. We optimize a Bi-ACT-style action chunks with an additional KL regularization term for the CVAE latent space, using AdamW for both the ResNet-18 backbone and the remaining modules and KL weight $\lambda_{\text{KL}}=1.0$.
At deployment, the trained policies run in real time on the physical robot using an NVIDIA RTX 3060 Ti GPU (8 GB VRAM).
Figure 12: Autonomous execution of Bi-AQUA under lighting conditions.
@@ IV-B3 Training Configuration
IV-C Experimental Results
IV-C1 Lighting Robustness and Ablation
Table III summarizes task success rates across all eight lighting conditions, including three unseen at training time (cyan, purple, changing). Fig. 12 shows the autonomous execution of Bi-AQUA. Bi-AQUA achieves consistently high robustness, attaining 100% success in seven of the eight lighting conditions and maintaining 80% success even under blue light, which is the most challenging condition due to severe wavelength-dependent attenuation. Crucially, Bi‑AQUA achieves 100% and clearly outperforms all baselines in the dynamic changing mode, where lighting cycles through all spectra every 2 sec. This setting induces rapid, intra-horizon appearance shifts that invalidate the stationarity assumptions of standard visuomotor policies.
@@ IV-C1 Lighting Robustness and Ablation
In contrast, the lighting-agnostic Bi-ACT baseline succeeds only under white (100%) and partially under red (20%), and fails completely in the remaining six conditions despite using the same demonstrations. This sharp performance gap demonstrates that underwater visuomotor control fundamentally requires explicit lighting modeling rather than relying on passive robustness of the visual backbone.
The ablations in Table II isolate the role of each lighting-aware component. Bi-ACT+LE-Token, which uses a Lighting Encoder to extract lighting embeddings and adds a lighting token to the transformer encoder input (conditioning the decoder via memory) without modifying the visual features via FiLM, fails in all lighting conditions. This indicates that high-level conditioning alone is insufficient for correcting color shifts and contrast changes introduced by underwater propagation.
Bi-ACT+LE-FiLM, which uses a Lighting Encoder to extract lighting embeddings and applies FiLM modulation to the final ResNet layer (layer4) but omits the lighting token, performs substantially better, achieving up to 100% success in five static or unseen lighting conditions. However, its success drops sharply to 20% in the dynamic changing mode, suggesting that FiLM alone cannot track rapid temporal lighting transitions.
Only the full Bi-AQUA model—combining Lighting Encoder, FiLM modulation, and a sequence-level lighting token—maintains uniformly high performance across all lighting conditions. This confirms that hierarchical lighting reasoning, spanning both feature modulation and sequence-level conditioning, is essential for underwater manipulation under severe lighting variability.
Fig. 13 further analyzes task execution time. Averaged over all lighting conditions, Bi-ACT takes $20.17$ sec per successful rollout, Bi-ACT+LE-FiLM $19.61$ sec, Bi-AQUA $15.73$ sec, and human teleoperation (Train) $15.39$ sec. Although Bi-ACT+LE-FiLM achieves high success in many static lighting settings, it consistently completes the task several seconds slower than Bi-AQUA and the teleoperator, reflecting cautious, stop-and-go behavior under difficult colors. Bi-AQUA, by contrast, closely matches the teleoperation reference in both mean time and variance, suggesting that explicit lighting modeling not only increases success rates but also preserves efficient, human-like trajectories instead of sacrificing speed for robustness. The per-lighting breakdown in Fig. 13 shows that this advantage holds across all lighting modes, including the dynamic changing condition, with Bi-AQUA maintaining consistently short and stable execution times.
Figure 13: Execution time profiles across the task stages under different lighting conditions.
IV-C2 Generalization to Novel Objects and Disturbances

### v2

@@ Abstract
Underwater robotic manipulation remains challenging because lighting variation, color attenuation, scattering, and reduced visibility can severely degrade visuomotor policies. We present Bi-AQUA, the first underwater bilateral control-based imitation learning framework for robot arms that explicitly models lighting within the policy. Bi-AQUA integrates transformer-based bilateral action chunking with a hierarchical lighting-aware design composed of a label-free Lighting Encoder, FiLM-based visual feature modulation, and a lighting token for action conditioning. This design enables adaptation to static and dynamically changing underwater illumination while preserving the force-sensitive advantages of bilateral control, which are particularly important in long-horizon and contact-rich manipulation. Real-world experiments on underwater pick-and-place, drawer closing, and peg extraction tasks show that Bi-AQUA outperforms a bilateral baseline without lighting modeling and achieves robust performance under seen, unseen, and changing lighting conditions. These results highlight the importance of combining explicit lighting modeling with force-aware bilateral imitation learning for reliable underwater manipulation. For additional material, please check: https://mertcookimg.github.io/bi-aqua
@@ I Introduction
Underwater robotic manipulation is highly challenging because scene appearance can change drastically within seconds due to shifts in underwater lighting spectrum, intensity, and direction [1]. Wavelength-dependent attenuation, scattering, turbidity, and bubbles disrupt visual consistency, often causing severe action drift in standard visuomotor policies. Although underwater image enhancement methods improve perceptual quality [2, 3], they do not address the core control problem: closed-loop policies must adapt both perception and action generation to rapidly changing illumination. While traditional underwater manipulation relies on teleoperation, recent work explores RL and IL-based autonomy [4, 5, 6]. However, existing approaches either depend purely on vision or employ unilateral control without force feedback, and none explicitly model lighting as a latent factor within the visuomotor policy.
@@
Figure 2: Unilateral Control-based Imitation Learning
@@
II Related Work
II-A Underwater Vision and Manipulation
Underwater perception is degraded by wavelength-dependent attenuation, scattering, turbidity, and non-uniform illumination [1]. Extensive research addresses image restoration through physics-based, learning-based, and hybrid approaches [3, 26, 27], as well as large-scale datasets and low-light correction [2]. However, these methods optimize image quality rather than closed-loop manipulation performance.
Autonomous underwater manipulation has evolved from teleoperation [21] toward RL and IL-based control for navigation and manipulation [4, 5, 6]. Nevertheless, existing systems either rely purely on vision or lack integration of force feedback and explicit lighting modeling, limiting robustness under dynamic underwater illumination.
Therefore, Bi-AQUA shifts the focus from image enhancement to manipulation performance, integrating force-sensitive bilateral control with explicit lighting modeling for robust underwater operation.
@@ II-B Bilateral Control-based Imitation Learning
Unilateral teleoperation systems such as ALOHA and Mobile ALOHA [8, 9, 10] depend solely on visual feedback, which restricts robustness in contact-rich tasks. Bilateral control-based imitation learning (Bi-IL) addresses this limitation by exchanging both position and force signals between leader and follower robots [7, 17, 24].
Recent works integrate images into Bi-IL, notably Bi-ACT [12, 13], but evaluations are limited to visually stable terrestrial environments. Underwater settings introduce severe photometric distortions, rendering proprioception-only approaches perceptually insufficient and vision-based Bi-IL vulnerable to lighting shifts. To our knowledge, no prior work unifies Bi-IL, real underwater deployment, and explicit lighting modeling within a single visuomotor policy.
Bi-AQUA extends bilateral imitation learning to real underwater environments by explicitly incorporating lighting-aware visual conditioning into a force-sensitive control framework.
@@ II-C Robust Visual Policies and Lighting Modeling
Visuomotor policies are highly sensitive to photometric shifts. Robustness is commonly improved through domain randomization [28] or physically based relighting [19], but these approaches treat illumination implicitly and do not provide an internal lighting representation within the policy.
Conditional modulation methods such as FiLM [29] enable feature adaptation [30], yet lighting is typically not disentangled as a structured latent factor influencing both perception and action generation.
In contrast, Bi-AQUA encodes lighting as a label-free latent variable and integrates it hierarchically via FiLM modulation and a dedicated lighting token, enabling adaptation to rapidly changing underwater illumination within the control policy itself.
Taken together, prior work has addressed underwater perception, bilateral imitation learning, and visual robustness largely in isolation. Bi-AQUA bridges these directions by unifying force-sensitive bilateral control with explicit, hierarchical lighting modeling within a single visuomotor policy for real underwater manipulation.
@@ III-B Data Collection
where $\theta$ and $\tau$ denote joint position and torques, respectively. Position is obtained from encoders, velocities by differentiation, and torques by a disturbance observer (DOB) [31] and a reaction force observer (RFOB) [32], without dedicated force sensors.
@@ IV-A Experimental Setup
We evaluate Bi-AQUA on a bilateral leader–follower system shown in Fig. 8. Each arm has three revolute joints (J1–J3) and a gripper (J4), actuated by Dynamixel XW540-T260 (IP68) servo motors. The follower arm is mounted on a 200 mm pedestal inside a water tank of $900\times 450\times 450$ $m$.
@@
To create reproducible underwater lighting variation, we install an RGBW LED aquarium light (FEDOUR RD10-400RGBW, 38 cm, 7 W, 21 LEDs, IP68) beneath the follower pedestal and use eight lighting modes: red, blue, green, white, rgbw, cyan, purple, and changing. All static modes correspond to the manufacturer’s 100% brightness presets, while the dynamic changing mode uses the built-in “M5” program, which cycles through colors every 2 s. This setup enables controlled evaluation under both static and rapidly varying illumination.
@@
Bi-AQUA receives RGB observations from two cameras (TIER IV C1 120): a gripper-mounted camera for local close-up views and a top camera observing the global workspace.
IV-B Tasks and Evaluation Protocol
IV-B1 Common Protocol
Across all tasks, we use a unified train/test lighting protocol to isolate the effect of lighting variation. Unless otherwise noted, we collect bilateral teleoperation demonstrations under five training lighting modes (red, blue, green, white, rgbw) and evaluate under all eight modes, including two unseen colors (cyan, purple) and the dynamic changing condition. Each test configuration is evaluated over 5 autonomous rollouts, and we report success rate (%).
We evaluate three manipulation tasks with progressively increasing temporal horizon and contact complexity:
  • Pick-and-Place Task: a lighting-robustness benchmark in which the robot transports an object across the tank.
  • Drawer Closing Task: a long-horizon sequential task with an additional drawer-closing operation.
  • Peg Extraction Task: a contact-rich precision manipulation task requiring stable grasping and controlled disengagement.
IV-B2 Pick-and-Place Task
Figure 9: Overview of underwater manipulation tasks
@@ IV-B2 Pick-and-Place Task
Pick-and-Place task is designed to directly stress lighting robustness in underwater manipulation policy. The follower must grasp the polyurethane block on the right side of the tank and place it into a basket on the opposite side, passing through four stages: #0 initial pose, #1 pick up, #2 transport, #3 place. A rollout is successful if the object ends inside the basket without being dropped.
To test robustness beyond the training object and scene, we additionally evaluate (a) a black rubber block, (b) a blue sponge, and (c) bubbles around the polyurethane block (Fig. 9). These settings are evaluated under the same eight lighting modes with the same success criterion.
IV-B3 Drawer Closing Task
Drawer closing task extends Pick-and-Place task into a long-horizon sequential task by appending a drawer closing operation. After placing the object into the target container, the robot approaches a drawer and performs a closing contact-rich motion. We define full-task success if both conditions hold: (i) the object is placed in the target container and (ii) the drawer reaches the closed state.
IV-B4 Peg Extraction Task
Peg Extraction evaluates contact-rich precision manipulation under severe underwater photometric variation. At the beginning of each rollout, the peg is partially inserted into a mating hole. The robot must establish a stable grasp and extract the peg without losing contact stability. The radial clearance between the peg and the hole is 1.5 mm, resulting in tight geometric tolerance and substantial frictional contact along the insertion surface. Consequently, successful extraction requires careful modulation of grasp force and pulling motion to overcome friction while avoiding excessive force that may destabilize the grasp. A rollout is considered successful if the peg is completely removed from the hole and securely held by the robot at the end of the episode.
Figure 10: Bilateral teleoperation data collection under five training lighting modes.
IV-C Dataset and Training Configuration
IV-C1 Demonstration Collection and Preprocessing
We use the bilateral control data-collection pipeline for all tasks. The policy is trained and evaluated only from follower-side observations, while leader joint states are used as action supervision.
For each task, we collect 10 bilateral teleoperation episodes (two per training lighting mode) under five training conditions: red, blue, green, white, and rgbw. For Pick-and-Place, demonstrations follow a consistent motion pattern so that the main variation comes from lighting rather than operator strategy.
@@ IV-C1 Demonstration Collection and Preprocessing
Two RGB cameras (top and gripper views) record at 30 Hz with resolution $1920\times 1280$, and images are resized to $480\times 320$. Joint position, velocity, and torque for $N=4$ joints are recorded from both arms at 1,000 Hz. For training, joint data are downsampled to 100 Hz, and images are aligned to the same rate by selecting the nearest frame at each 10 ms timestamp. Each training sample therefore consists of two RGB views, the follower joint state, and the corresponding leader joint state.
@@
Although demonstrations are collected under known lighting modes, no lighting IDs or manual labels are provided. The Lighting Encoder is trained end-to-end through the imitation objective and learns a latent representation directly from raw images.
IV-C2 Compared Methods and Ablations (Pick-and-Place task)
TABLE II: Ablation of lighting-aware components.
@@ IV-C2 Compared Methods and Ablations (Pick-and-Place task)
All variants share the same Bi-ACT-style visuomotor backbone and differ only in three lighting-aware components: (i) a 64-D Lighting Encoder, (ii) FiLM modulation in the visual backbone, and (iii) a lighting token added to the transformer encoder input. Baseline and ablation comparisons are conducted on Pick-and-Place to isolate the effect of explicit lighting modeling, while Drawer Closing and Peg Extraction are used to evaluate task generality for Bi-AQUA.
IV-C3 Training Configuration
All models are trained offline with the same preprocessing and optimization settings. We use a Bi-ACT-style CVAE action-chunking policy with KL regularization, optimized by AdamW with learning rate $1\times 10^{-5}$ for both the ResNet-18 backbone and the remaining modules. The KL weight is set to $\lambda_{\text{KL}}=1.0$. The transformer uses 8 attention heads, hidden dimension 512, and feedforward dimension 3,200.
Figure 11: Execution of Bi-AQUA under lighting conditions.
IV-D Experimental Results
IV-D1 Pick-and-Place task: Lighting Robustness and Ablation
@@
Figure 12: t-SNE of each camera.
Figure 13: Execution time profiles (Pick-and-Place task).
TABLE IV: Success rate (%) across lighting and multi-stage task
Method Trained Lighting Untrained Lighting
red blue green white rgbw cyan purple changing
Multi-stage drawer closing contact-rich task.
Bi-AQUA (w/o Force) 40 40 100 60 40 100 100 40
Bi-AQUA (Ours) 100 100 100 80 80 100 80 80
Peg-out-Hole contact-rich task.
Bi-AQUA (w/o Force) 0 40 60 60 80 40 0 0
Bi-AQUA (Ours) 100 100 100 100 100 100 80 100
Table III reports Pick-and-Place success rates under all eight lighting modes, including two unseen colors (cyan and purple) and the dynamic changing condition. Bi-AQUA shows strong robustness, achieving 100% success in seven of the eight modes and maintaining 80% under blue illumination, which is particularly challenging because of wavelength-dependent attenuation and reduced underwater contrast. Notably, Bi-AQUA also achieves 100% success in the changing mode, where the illumination switches every 2 s and scene appearance changes substantially within a single action horizon. These results indicate robustness to both static photometric shifts and temporally varying illumination.
@@
In contrast, the lighting-agnostic Bi-ACT baseline succeeds only under white illumination (100%) and partially under red illumination (20%), while failing in all remaining conditions despite using the same demonstrations. This gap shows that explicit lighting modeling is essential for reliable underwater visuomotor control.
The ablation results in Table II further clarify the role of each component. Bi-ACT+LE-Token fails across all modes, indicating that sequence-level conditioning alone cannot compensate for underwater color and contrast distortion. Bi-ACT+LE-FiLM performs strongly in several static and unseen lighting conditions, confirming the importance of lighting-aware visual feature modulation, but drops sharply in the dynamic changing mode. Only the full Bi-AQUA model, combining the Lighting Encoder, FiLM modulation, and the lighting token, maintains uniformly high success across all conditions, demonstrating that hierarchical lighting adaptation is necessary for robust underwater manipulation.
This interpretation is supported by the t-SNE visualization in Fig. 12. The learned lighting embeddings form meaningful clusters by lighting condition, especially for the gripper camera and the fused multi-view representation. The top-camera embedding shows larger overlap, likely due to water-surface reflection and global scene ambiguity, whereas averaging the two camera embeddings yields clearer separation, including for unseen lighting conditions. These results suggest that the Lighting Encoder learns a structured latent representation of underwater illumination without explicit labels, and that multi-view fusion improves lighting discriminability.
Fig. 13 shows the execution time analysis. Averaged over all lighting conditions, Bi-AQUA completes the task in $15.73$ s, closely matching human bilateral teleoperation at $15.39$ s, while the baselines are slower (Bi-ACT: $20.17$ s; Bi-ACT+LE-FiLM: $19.61$ s). Thus, the proposed lighting-aware design improves robustness without sacrificing execution efficiency, enabling smooth and efficient task execution under severe underwater photometric variation.
IV-D2 Pick-and-Place task: Generalization to Novel Objects and Disturbances